# USV-AUV Collaboration Framework for Underwater Tasks under Extreme Sea Conditions

### v1

@@ Abstract
Autonomous underwater vehicles (AUVs) are valuable for ocean exploration due to their flexibility and ability to carry communication and detection units. Nevertheless, AUVs alone often face challenges in harsh and extreme sea conditions. This study introduces a unmanned surface vehicle (USV)–AUV collaboration framework, which includes high-precision multi-AUV positioning using USV path planning via Fisher information matrix optimization and reinforcement learning for multi-AUV cooperative tasks. Applied to a multi-AUV underwater data collection task scenario, extensive simulations validate the framework’s feasibility and superior performance, highlighting exceptional coordination and robustness under extreme sea conditions. The simulation code will be made available as open-source to foster future research in this area.
@@ I Introduction
To address the aforementioned challenges, numerous researchers have concentrated on the development of USV-AUV co-localization and intelligent multi-AUV collaboration. Bahr et al. introduced a distributed algorithm that leverages multiple AUVs to dynamically determine the locally optimal position of the beacon vehicle, utilizing data from the survey vehicle’s broadcast communication [11]. Vasilijevi et al. constructed the Internet of underwater things using unmanned surface vehicles (USVs) to enhance underwater positioning efficiency [12]. Jiang et al. applied a multi-agent proximal policy optimization reinforcement learning (RL) algorithm to direct efficient and energy-saving data collection for AUV swarms in unknown environments based on an objective uncertainty map [13]. Wang et al. proposed a collaborative data collection strategy for multi-AUVs employing local-global deep Q-learning and data value, classifying data into urgent and non-urgent categories to facilitate hybrid data collection meeting various temporal demands [14]. Nevertheless, these methods encounter limitations in complex underwater environments: as the number of AUVs and sensors increases, the underwater acoustic channel becomes more intricate, computational complexity rises [15], and there are stringent requirements for battery life and operational costs of the equipment. Furthermore, conventional AUV swarm control techniques such as heuristic algorithms [16], neural networks [7], and game theory [17], although effective for specific tasks, heavily rely on extensive prior information. In the absence of such information, the performance of these methods significantly declines, particularly under extreme sea conditions.
@@ III-B Experiment Results and Analysis
Moreover, we utilized the expert policy trained with the DDPG algorithm to guide the multi-AUV system in an underwater data collection task under extreme sea conditions. The trajectories of the AUVs and the USV during a single RL training episode are depicted in Fig. (4a). To further evaluate the advantages of the USV-AUV collaboration framework, we also presented the positioning error of the multi-AUV system. We examined three different scenarios: employing USV path planning based on FIM optimization, fixing the USV at coordinates (0,0), and fixing it at (100,100). As shown in Fig. (4b), the first approach achieves the lowest positioning error, illustrating the superior performance of USBL positioning via USV path planning using FIM optimization, even under extreme sea conditions.
@@ IV Conclusion
In this paper, we propose a USV–AUV collaboration framework. The two parts of our framework, including high-precision multi-AUV location using USV path planning by FIM optimization and RL training for multi-AUV cooperative tasks, collaboratively enhance the performance of multi-AUV underwater tasks in extreme sea conditions. Experiment results from underwater data collection task verify the superior feasibility and performance of the proposed framework, which effectively embodies high coordination between USV and AUV while showcasing excellent robustness in extreme sea conditions. To accelerate the relevant research in this field, the code for simulation will be released as open-source in the future.

### v2

@@ Abstract
Autonomous underwater vehicles (AUVs) are valuable for ocean exploration due to their flexibility and ability to carry communication and detection units. Nevertheless, AUVs alone often face challenges in harsh and extreme sea conditions. This study introduces a unmanned surface vehicle (USV)–AUV collaboration framework, which includes high-precision multi-AUV positioning using USV path planning via Fisher information matrix optimization and reinforcement learning for multi-AUV cooperative tasks. Applied to a multi-AUV underwater data collection task scenario, extensive simulations validate the framework’s feasibility and superior performance, highlighting exceptional coordination and robustness under extreme sea conditions. To accelerate relevant research in this field, we have made the simulation code available as open-source1.
@@ I Introduction
To address the aforementioned challenges, numerous researchers have concentrated on the development of USV-AUV co-localization and intelligent multi-AUV collaboration. Bahr et al. introduced a distributed algorithm that leverages multi-AUV to dynamically determine the locally optimal position of the beacon vehicle, utilizing data from the survey vehicle’s broadcast communication [11]. Vasilijevi et al. constructed the Internet of underwater things using unmanned surface vehicles (USVs) to enhance the underwater positioning efficiency [12]. Jiang et al. applied a multi-agent proximal policy optimization reinforcement learning (RL) algorithm to direct efficient and energy-saving data collection for AUVs in unknown environments based on an objective uncertainty map [13]. Wang et al. proposed a collaborative data collection strategy for multi-AUV employing local-global deep Q-learning and data value, to facilitate hybrid data collection meeting various temporal demands [14]. Nevertheless, these methods encounter limitations in complex environments: as number of AUVs and sensors increases, the underwater acoustic channel becomes more intricate, computational complexity rises [15], and there are stringent requirements for battery and operational costs of the equipment. Furthermore, traditional AUV swarm control techniques such as heuristic algorithms [16], neural networks [7], and game theory [17], although effective for specific tasks, heavily rely on extensive prior information. In the absence of such information, performance of these methods significantly declines, particularly under extreme sea conditions.
@@ III-B Experiment Results and Analysis
Moreover, we utilized the expert policy trained with the DDPG algorithm to guide the multi-AUV system in an underwater data collection task under extreme sea conditions. The trajectories of the AUVs and the USV during a single RL training episode are depicted in Fig. 4(a). To further evaluate the advantages of the USV-AUV collaboration framework, we also presented the positioning error of the multi-AUV system. We examined three different scenarios: employing USV path planning based on FIM optimization, fixing the USV at coordinates (0,0), and fixing it at (100,100). As shown in Fig. 4(b), the first approach achieves the lowest positioning error, illustrating the superior performance of USBL positioning via USV path planning using FIM optimization, even under extreme sea conditions.
@@ IV Conclusion
In this paper, we propose a USV–AUV collaboration framework. The two parts of our framework, including high-precision multi-AUV location using USV path planning by FIM optimization and RL training for multi-AUV cooperative tasks, collaboratively enhance the performance of multi-AUV underwater tasks in extreme sea conditions. Experiment results from underwater data collection task verify the superior feasibility and performance of the proposed framework, which effectively embodies high coordination between USV and AUV while showcasing excellent robustness in extreme sea conditions. To accelerate relevant research in this field, we have made the simulation codes available as open-source at https://github.com/360ZMEM/USV-AUV-colab.